OpenStudy (anonymous):

If triangle abc and triangle pqr are similar triangles find qr and pr? AB=15, AC=20, BC=25, PQ=21

OpenStudy (anonymous):

PQ/AB=21/15=1.4

OpenStudy (anonymous):

PR/AC=PR/20=1.4 => PR=1.4*20

OpenStudy (anonymous):

QR=1.4*25

geerky42 (geerky42):

 Then use ratio proportion to find missing side length. Like what AJ01 did.